# Parallel AutoRegressive Models for Multi-Agent Combinatorial Optimization

### v2

@@ 6 Conclusion
We introduced PARCO, a learning model to tackle multi-agent combinatorial optimization problems efficiently via parallel autoregressive solution construction. By integrating transformer-based Communication Layers, a Multiple Pointer Mechanism, and Priority-based Conflict Resolution, PARCO enables effective agent coordination and significantly reduces computational latency. Our extensive experiments on multi-agent vehicle routing and scheduling demonstrate that PARCO consistently outperforms state-of-the-art learning-based solvers, achieving higher solution quality while maintaining superior efficiency.
@@
Although PARCO can efficiently solve multi-agent CO problems with a defined number of agents $M$, it cannot be directly applied to CO tasks where solutions are constructed via an unspecified $M$. In future work, we plan to explore multi-agent CO problems with an unspecified number of agents, which could be achieved by either rolling out a batch of several values of $M$ until an optimal solution is reached or by employing a prediction module to predict an optimal agent number.
@@ Acknowledgements
We thank people in the AI4CO open research community who have contributed to RL4CO, which this work is based on. This work was supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean government(MSIT)[2022-0-01032, Development of Collective Collaboration Intelligence Framework for Internet of Autonomous Things].
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 HCVRP
The min-max HCVRP (Heterogeneous Capacitated Vehicle Routing Problem) consists of ${M}$ agents sequentially visiting customers to satisfy their demands, with constraints including each customer can be visited exactly once and the amount of demand satisfied by a single vehicle in a trip cannot exceed its capacity, which can be reloaded by going back to the depot. The goal is to minimize the makespan, i.e., the worst route.
Consider a problem with $N+1$ nodes (including $N$ customers and a depot) and ${M}$ vehicles. The depot is indexed as 0, and customers are indexed from 1 to $N$.
@@
$N$ Number of customer nodes (excluding depot)
${M}$ Number of vehicles
@@
$x_{ijk}$$\left\{\begin{array}[]{ll}1&\text{if vehicle }k\text{ travels directly from node }i\text{ to node }j\\ 0&\text{otherwise}\end{array}\right.$
@@
Constraint Explanations: The formulation is subject to several constraints that define the feasible solution space. Equation (11) ensures that each customer is visited exactly once by one vehicle. The flow conservation constraint (12) guarantees that each vehicle that enters a node also leaves that node, maintaining route continuity. Demand satisfaction is enforced by constraint (13), which ensures that the difference in load before and after serving a customer equals the customer’s demand. The vehicle capacity constraint (14) ensures that the load carried by a vehicle does not exceed its capacity and is sufficient to meet the next customer’s demand.
@@ A.2 OMDCPDP
$N$ Number of pickup-delivery pairs
${M}$ Number of vehicles
@@
$x_{ijk}$$\left\{\begin{array}[]{ll}1&\text{if vehicle }k\text{ travels directly from node }i\text{ to node }j\\ 0&\text{otherwise}\end{array}\right.$
@@
Constraints Explanations: Equation (18) ensures that each node is visited exactly once. Constraint (19) guarantees that each vehicle starts from its designated depot. The flow conservation constraint (20) ensures route continuity for each vehicle. Equation (21) defines the relationship between x and y variables. Time consistency is enforced by constraint (22), while (23) ensures that pickups are visited before their corresponding deliveries. Constraints (24) and (25) manage the load changes during pickup and delivery operations. Finally, the vehicle capacity constraint (26) ensures that the load never exceeds the vehicle’s stacking limit.
@@ A.3 FFSP
We use the mathematical model outlined in Kwon et al. (2021) to define the FFSP:
@@
$X_{ijk}$$\left\{\begin{array}[]{ll}1&\text{if job }j\text{ is assigned to machine }k\text{ in stage }i\\ 0&\text{otherwise}\end{array}\right.$
$Y_{ilj}$$\left\{\begin{array}[]{ll}1&\text{if job }l\text{ is processed earlier than job }j\text{ in stage }i\\ 0&\text{otherwise}\end{array}\right.$
@@
Constraint Explanations: Here, the objective function Equation 29 minimizes the makespan of the resulting schedule, that is, the completetion time of the job that finishes last. The schedule has to adhere to several constraints: First, constraint set Equation 30 ensures that each job is assigned to exactly one machine at each stage. Constraint sets Equation 31 through Equation 35 define the precedence relationships between jobs within a stage. Specifically, constraint set Equation 31 ensures that a job has no precedence relationship with itself. Constraint set Equation 32 ensures that the total number of precedence relationships in a stage equals $N-m_{i}$ minus the number of machines with no jobs assigned. Constraint set Equation 33 dictates that precedence relationships can only exist among jobs assigned to the same machine. Additionally, constraint sets Equation 34 and Equation 35 restrict a job to having at most one preceding job and one following job.
Moving on, constraint set Equation 36 specifies that the completion time of a job in the first stage must be at least as long as its processing time in that stage. The relationship between the completion times of a job in consecutive stages is described by constraint set Equation 37. Finally, constraint set Equation 38 ensures that no more than one job can be processed on the same machine simultaneously.
@@ B.1.1 Baselines
We follow the experimental setup of Liu et al. (2024b) for baselines, with additional baselines hyperparameter details reported in their respective papers.
@@
The Slack Induction by String Removals (SISR) approach (Christiaens & Vanden Berghe, 2020) offers a heuristic method for addressing vehicle routing problems (VRPs), focusing on simplifying the optimization process. It combines techniques for route dismantling and reconstruction, along with vehicle fleet minimization strategies. SISR is applied across various VRP scenarios, including those with specific pickup and delivery tasks. In our experiments, we adhere to the hyperparameters provided in the original paper with $\bar{c}=10,L^{\mathrm{max}}=10,\alpha=10^{-3},\beta=10^{-2},T_{0}=100,T_{f}=1,\mathrm{iter}=3\times 10^{5}\times N$.
@@
The Genetic Algorithm (GA) (Karakatič & Podgorelec, 2015) is used to address vehicle routing problems (VRPs) and other NP-hard challenges by simulating natural evolutionary processes. GA generates adequate solutions with reasonable computational resources. Our experiment follows the same carefully tuned hyperparameters from (Liu et al., 2024b) with $n=200,\mathrm{iter}=40\times N,P_{m}=0.8,P_{c}=1$.
@@
The Simulated Annealing (SA) method (İlhan, 2021) targets the capacitated vehicle routing problem (CVRP) using a population-based approach combined with crossover operators. It incorporates local search and the improved 2-opt algorithm to refine routes alongside crossover techniques to speed up convergence. In our experiment, we follow the same carefully tuned hyperparameters from (Liu et al., 2024b) with $T_{0}=100,T_{f}=10^{-7},L=20\times N,\alpha=0.98$.
@@
The Attention Model (AM) (Kool et al., 2019a) applies the attention mechanism to tackle combinatorial optimization problems like the Traveling Salesman and Vehicle Routing Problems. It utilizes attention layers for model improvement and trains using REINFORCE with deterministic rollouts. In our studies, we adopt adjustments from the DRLLi framework, which involves selecting vehicles sequentially and then choosing the next node for each. Additionally, vehicle-specific features are incorporated into the context vector generation to distinguish between different vehicles.
@@
The Equity-Transformer (ET) approach (Son et al., 2024) addresses large-scale min-max routing problems by employing a sequential planning approach with sequence generators like the Transformer. It focuses on equitable workload distribution among multiple agents, applying this strategy to challenges like the min-max multi-agent traveling salesman and pickup and delivery problems. In our experiments, we modify the decoder mask in ET to generate feasible solutions for HCVRP and integrate vehicle features into both the input layer and the context encoder, similarly to the setting of Liu et al. (2024b).
@@
The Decoupling-Partition-Navigation (DPN) approach (Zheng et al., 2024a) is a SOTA sequential planning AR baseline that tackles min-max vehicle routing problems (min-max VRPs) by explicitly separating the tasks of customer partitioning and route navigation. It introduces a Partition-and-Navigation (P&N) Encoder to learn distinct embeddings for these tasks, an Agent-Permutation-Symmetric (APS) loss to leverage routing symmetries, and a Rotation-Based Positional Encoding to enhance generalization across different depot locations. We employ a similar setting as ET.
@@
The DRL approach for solving HCVRP by Li et al. (2022) employs a transformer architecture similar to Kool et al. (2019a) in which the vehicle and node selection happens in two steps via a two selection decoder, thus requiring two actions. We employ their original model with additional context of variable vehicle speeds, noting that in the original setting each model was trained on a single distribution of number of agents ${M}$, each with always the same characteristics.
@@
The 2D Array Pointer network (2D-Ptr) (Liu et al., 2024b) addresses the heterogeneous capacitated vehicle routing problem (HCVRP) by using a dual-encoder setup to map vehicles and customer nodes effectively. This approach facilitates dynamic, real-time decision-making for route optimization. Its decoder employs a 2D array pointer for action selection, prioritizing actions over vehicles. The model is designed to adapt to vehicle and customer numbers changes, ensuring robust performance across different scenarios.
@@ B.1.2 Datasets
Neural baselines were trained with the specific number of nodes $N$ and number of agents ${M}$ they were tested on. In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $N\sim\mathcal{U}(60,100)$ and $m\sim\mathcal{U}(3,7)$. As we show in Table 1, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution. The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space. The depot location is similarly sampled using the same uniform distribution. The demand $d_{i}$ for each customer $i$ is also drawn from a uniform distribution $\mathcal{U}(1,10)$, with the depot having no demand, i.e., $d_{0}=0$. Each vehicle $m$, where $m=1,\ldots,{M}$, is assigned a capacity $Q_{m}$ sampled from a uniform distribution $\mathcal{U}(20,41)$. The speed $f_{m}$ of each vehicle is uniformly distributed within the range $\mathcal{U}(0.5,1.0)$.
@@
Testing is performed on the 1280 instances per $(N,M)$ test setting from Liu et al. (2024b). In Table 1, (g.) refers to the greedy performance of the model, i.e., taking a single trajectory by taking the maximum action probability; (s.) refers to sampling $1280$ solutions in the latent space and selecting the one with the lowest cost (i.e., highest reward).
@@ B.1.3 PARCO Network Hyperparameters
Initial Embedding. This layer projects initial raw features to hidden space. For the depot, the initial embedding is the positional encoding of the depot’s location $X_{0}$. For agents, the initial embedding is the encoding for the initial location, capacity, and speed. Main Encoder. we employ $L=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang & Sennrich, 2019) as normalization before the MHA and the MLP.
@@
We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang & Sennrich, 2019) as normalization before the MHA and the MLP. Unlike the encoder layer, which acts between all $M+N$ problem tokens, communication layers are lighter because they communicate between ${M}$ agents.
@@
We use the Priority-based Conflict Handler guided by the model output probability for managing conflicts: priority is given to the agent whose probability of selecting the conflicting action is the highest (see Section 4.6).
@@ B.1.4 PARCO Training Hyperparameters
Unlike baselines, which are trained and tested on the same distribution, we train a single PARCO model that can effectively generalize over multiple size and agent distributions thanks to our flexible structure. We train PARCO with RL via SymNCO (Kim et al., 2022) with $K=10$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer (Kingma & Ba, 2014) with a total batch size $512$ (using 4 GPUs in Distributed Data Parallel configuration) and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs. For each epoch, we sample $4\times 10^{5}$ randomly generated data. Training takes around 15 hours in our configuration.
@@ B.2 OMDCPDP
The setting introduced in OMDCPDP is a more general and realistic setting than the one introduced in Zong et al. (2022), particularly due to the multiple depots and the global lateness objective function which is harder to optimize than vanilla min-sum.

### v3

@@ 6 Conclusion
We introduced PARCO, a learning model to tackle multi-agent combinatorial optimization problems efficiently via parallel autoregressive solution construction. By integrating transformer-based Communication Layers, a Multiple Pointer Mechanism, and Priority-based Conflict Resolution, PARCO enables effective agent coordination and significantly reduces computational latency. Our extensive experiments on multi-agent vehicle routing and scheduling demonstrate that PARCO consistently outperforms state-of-the-art learning-based solvers with better solution quality and higher efficiency.
@@
Although PARCO can efficiently solve multi-agent CO problems with a defined number of agents $M$, it cannot be directly applied to CO tasks where solutions are constructed via an unspecified $M$. In future work, we plan to explore multi-agent CO problems with an unspecified number of agents, which could be achieved by either rolling out a batch of several values of $M$ until an optimal solution is reached or by employing a prediction module to predict an optimal agent number. We defer additional discussions to §˜C.1.
@@ Acknowledgements
We are deeply grateful to the members of the AI4CO open research community for their invaluable contributions to PARCO and related projects, including RL4CO. We also thank the anonymous reviewers who greatly helped improve our paper with their constructive feedback. This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant, funded by the Korean government (MSIT) [Grant No. 2022-0-01032, Development of Collective Collaboration Intelligence Framework for Internet of Autonomous Things]; National Research Foundation of Korea(NRF) grants funded by the Korea government(MSIT) (No. RS-2024-00410082 and No. RS-2025-00563763), and by the InnoCORE program of the Ministry of Science and ICT(N10250154).
Supplementary Material
@@ A.1 HCVRP
The min-max HCVRP (Heterogeneous Capacitated Vehicle Routing Problem) consists of M{M} agents sequentially visiting customers to satisfy their demands, with constraints including each customer can be visited exactly once and the amount of demand satisfied by a single vehicle in a trip cannot exceed its capacity, which can be reloaded by going back to the depot. The goal is to minimize the makespan, i.e., the worst route.
Consider a problem with $N+1$ nodes (including $N$ customers and a depot) and M{M} vehicles. The depot is indexed as 0, and customers are indexed from 1 to $N$.
@@
NN Number of customer nodes (excluding depot)
M{M} Number of vehicles
@@
[…]
@@
Constraint Explanations: The formulation is subject to several constraints that define the feasible solution space. Eq.˜10 ensures that each customer is visited exactly once by one vehicle. The flow conservation constraint (11) guarantees that each vehicle that enters a node also leaves that node, maintaining route continuity. Demand satisfaction is enforced by constraint (12), which ensures that the difference in load before and after serving a customer equals the customer’s demand. The vehicle capacity constraint (13) ensures that the load carried by a vehicle does not exceed its capacity and is sufficient to meet the next customer’s demand.
@@ A.2 OMDCPDP
NN Number of pickup-delivery pairs
M{M} Number of vehicles
@@
[…]
@@
Constraints Explanations: Eq.˜17 ensures that each node is visited exactly once. Constraint (18) guarantees that each vehicle starts from its designated depot. The flow conservation constraint (19) ensures route continuity for each vehicle. Eq.˜20 defines the relationship between x and y variables. Time consistency is enforced by constraint (21), while (22) ensures that pickups are visited before their corresponding deliveries. Constraints (23) and (24) manage the load changes during pickup and delivery operations. Finally, the vehicle capacity constraint (25) ensures that the load never exceeds the vehicle’s stacking limit.
Visualization
We provide a visualization of a large-scale instance in Fig.˜5.
Figure 5: Real-world instance for the OMDCPDP problem in Seoul City, South Korea, with $N=1000$ locations and $m=100$ agents ( ) showing relations () of pickups ( ) and their respective deliveries ( ).
@@ A.3 FFSP
We use the mathematical model outlined in Kwon et al. [50] to define the FFSP:
@@
[…]
@@
Constraint Explanations: Here, the objective function Eq.˜28 minimizes the makespan of the resulting schedule, that is, the completion time of the job that finishes last. The schedule has to adhere to several constraints: First, constraint set (29) ensures that each job is assigned to exactly one machine at each stage. Constraint sets (30) through (34) define the precedence relationships between jobs within a stage. Specifically, constraint set (30) ensures that a job has no precedence relationship with itself. Constraint set (31) ensures that the total number of precedence relationships in a stage equals $N-m_{i}$ minus the number of machines with no jobs assigned. Constraint set (32) dictates that precedence relationships can only exist among jobs assigned to the same machine. Additionally, constraint sets (33) and (34) restrict a job to having at most one preceding job and one following job.
Moving on, constraint set (35) specifies that the completion time of a job in the first stage must be at least as long as its processing time in that stage. The relationship between the completion times of a job in consecutive stages is described by constraint set (36). Finally, constraint set (37) ensures that no more than one job can be processed on the same machine simultaneously.
@@ B.1.1 Baselines
We follow the experimental setup of Liu et al. [63] for baselines, with additional baselines hyperparameter details reported in their respective papers.
@@
The Slack Induction by String Removals (SISR) approach [12] offers a heuristic method for addressing vehicle routing problems (VRPs), focusing on simplifying the optimization process. It combines techniques for route dismantling and reconstruction, along with vehicle fleet minimization strategies. SISR is applied across various VRP scenarios, including those with specific pickup and delivery tasks. In our experiments, we adhere to the hyperparameters provided in the original paper with $\bar{c}=10,L^{\mathrm{max}}=10,\alpha=10^{-3},\beta=10^{-2},T_{0}=100,T_{f}=1,\mathrm{iter}=3\times 10^{5}\times N$.
@@
The Genetic Algorithm (GA) [39] is used to address vehicle routing problems (VRPs) and other NP-hard challenges by simulating natural evolutionary processes. GA generates adequate solutions with reasonable computational resources. Our experiment follows the same carefully tuned hyperparameters from [63] with $n=200,\mathrm{iter}=40\times N,P_{m}=0.8,P_{c}=1$.
@@
The Simulated Annealing (SA) method [32] targets the capacitated vehicle routing problem (CVRP) using a population-based approach combined with crossover operators. It incorporates local search and the improved 2-opt algorithm to refine routes alongside crossover techniques to speed up convergence. In our experiment, we follow the same carefully tuned hyperparameters from [63] with $T_{0}=100,T_{f}=10^{-7},L=20\times N,\alpha=0.98$.
@@
The Attention Model (AM) [47] applies the attention mechanism to tackle combinatorial optimization problems like the Traveling Salesman and Vehicle Routing Problems. It utilizes attention layers for model improvement and trains using REINFORCE with deterministic rollouts. In our studies, we adopt adjustments from the DRLLi framework, which involves selecting vehicles sequentially and then choosing the next node for each. Additionally, vehicle-specific features are incorporated into the context vector generation to distinguish between different vehicles.
@@
The Equity-Transformer (ET) approach [84] addresses large-scale min-max routing problems by employing a sequential planning approach with sequence generators like the Transformer. It focuses on equitable workload distribution among multiple agents, applying this strategy to challenges like the min-max multi-agent traveling salesman and pickup and delivery problems. In our experiments, we modify the decoder mask in ET to generate feasible solutions for HCVRP and integrate vehicle features into both the input layer and the context encoder, similarly to the setting of Liu et al. [63].
@@
The Decoupling-Partition-Navigation (DPN) approach [104] is a SOTA sequential planning AR baseline that tackles min-max vehicle routing problems (min-max VRPs) by explicitly separating the tasks of customer partitioning and route navigation. It introduces a Partition-and-Navigation (P&N) Encoder to learn distinct embeddings for these tasks, an Agent-Permutation-Symmetric (APS) loss to leverage routing symmetries, and a Rotation-Based Positional Encoding to enhance generalization across different depot locations. We employ a similar setting as ET.
@@
The DRL approach for solving HCVRP by Li et al. [55] employs a transformer architecture similar to Kool et al. [47] in which the vehicle and node selection happens in two steps via a two selection decoder, thus requiring two actions. We employ their original model with additional context of variable vehicle speeds, noting that in the original setting each model was trained on a single distribution of number of agents M{M}, each with always the same characteristics.
@@
The 2D Array Pointer network (2D-Ptr) [63] addresses the heterogeneous capacitated vehicle routing problem (HCVRP) by using a dual-encoder setup to map vehicles and customer nodes effectively. This approach facilitates dynamic, real-time decision-making for route optimization. Its decoder employs a 2D array pointer for action selection, prioritizing actions over vehicles. The model is designed to adapt to vehicle and customer numbers changes, ensuring robust performance across different scenarios.
@@ B.1.2 Datasets
Neural baselines were trained with the specific number of nodes $N$ and number of agents M{M} they were tested on. In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $N\sim\mathcal{U}(60,100)$ and $m\sim\mathcal{U}(3,7)$. As we show in Table˜1, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution. The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space. The depot location is similarly sampled using the same uniform distribution. The demand $d_{i}$ for each customer $i$ is also drawn from a uniform distribution $\mathcal{U}(1,10)$, with the depot having no demand, i.e., $d_{0}=0$. Each vehicle $m$, where $m=1,\ldots,{M}$, is assigned a capacity $Q_{m}$ sampled from a uniform distribution $\mathcal{U}(20,41)$. The speed $f_{m}$ of each vehicle is uniformly distributed within the range $\mathcal{U}(0.5,1.0)$.
@@
Testing is performed on the 1280 instances per $(N,M)$ test setting from Liu et al. [63]. In Table˜1, (g.) refers to the greedy performance of the model, i.e., taking a single trajectory by taking the maximum action probability; (s.) refers to sampling $1280$ solutions in the latent space and selecting the one with the lowest cost (i.e., highest reward).
@@ B.1.3 PARCO Network Hyperparameters
Initial Embedding. This layer projects initial raw features to hidden space. For the depot, the initial embedding is the positional encoding of the depot’s location $X_{0}$. For agents, the initial embedding is the encoding for the initial location, capacity, and speed. Main Encoder. we employ $L=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm [100] as normalization before the MHA and the MLP.
@@
We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm [100] as normalization before the MHA and the MLP. Unlike the encoder layer, which acts between all $M+N$ problem tokens, communication layers are lighter because they communicate between M{M} agents.
@@
We use the Priority-based Conflict Handler guided by the model output probability for managing conflicts: priority is given to the agent whose probability of selecting the conflicting action is the highest (see §˜4.6).
@@ B.1.4 PARCO Training Hyperparameters
Unlike baselines, which are trained and tested on the same distribution, we train a single PARCO model that can effectively generalize over multiple size and agent distributions thanks to our flexible structure. We train PARCO with RL via SymNCO [43] with $K=10$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer [46] with a total batch size $512$ (using 4 GPUs in Distributed Data Parallel configuration) and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs. For each epoch, we sample $4\times 10^{5}$ randomly generated data. Training takes around 15 hours in our configuration.
@@ B.2 OMDCPDP
The setting introduced in OMDCPDP is a more general and realistic setting than the one introduced in Zong et al. [108], particularly due to the multiple depots and the global lateness objective function which is harder to optimize than vanilla min-sum.